How many grams is 3.5 x 1024 molecules of P2O4?


1
Expert's answer
2021-04-07T11:31:58-0400

Solution:

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Moles of P2O4 = (3.5×1024 molecules of P2O4) × (1 mol / 6.022×1023 molecules) = 5.812 moles P2O4


The molar mass of P2O4 is 125.94 g/mol.

Hence,

Mass of P2O4 = (5.812 mol P2O4) × (125.94 g P2O4 / 1 mol P2O4) = 731.96 g P2O4 = 732.0 g P2O4

Mass of P2O4 = 732.0 g


Answer: 732.0 grams is 3.5×1024 molecules of P2O4.